OpenStudy (anonymous):

Rationalize and Simplify 5i - 2i ----- 1 - 2i

OpenStudy (anonymous):

multiply by 1+2i to realize the denominator you then get (3i - 6)/5

OpenStudy (anonymous):

wait what??can you explain?

OpenStudy (anonymous):

Your numerator = 5i -2i = 3i Now, your denominator has to be free of all i's. How do we cancel out the i in the denominator? We multiply 1-2i by 1+2i (but to be valid, we need to multiply the top by 1+2i as well) 3i(1+2i) = 3i +6i^2 = 3i - 6 (1+2i)(1-2i) = 1 -4i^2 = 1+4 = 5 3i-6/5